n an expression 2[2+2{39-2(17+2)}], which part has to be simplified first


1
Expert's answer
2021-05-07T11:31:47-0400

The given expression is-


2[2+2{39-2(17+2)}]


The priority of the operator are- /, *, +, -


So On simplification It gives-


"=2[2+2\times 1]\\

 =2(4)\\=8"